OpenStudy (danielletries):

Help! Physics problem! A radioactive sample contains 1.30 ✕ 1011 atoms of _47^108text(Ag) (half-life = 2.42 min) at some instant. Calculate the decay constant and the activity of the sample in mCi.

OpenStudy (yagosik):

Well, decay constant is easy:\[\tau=\frac{ T _{1/2} }{ \ln 2 }\approx 1,44T _{1/2}\]Activity is more problematic:\[A(t)=\frac{ \ln 2 }{ T _{1/2} }N _{0}2^{-t/T _{1/2}}\]